OpenStudy (anonymous):

A store sells T-shirts for $10 each and jeans for $20 a pair. Anna spends $150 on T-shirts and jeans. She buys x T-shirts and y pairs of jeans. The equation that represents the situation is . If Anna decides to buy only T-shirts and no jeans, she can buy T-shirts with the money she has.

OpenStudy (anonymous):

x+y=150 multiply by the cost for each solve x $10x+$20y=$150 10x+20y-20y= 150 -20y 10x=-20y +150 10x/10=-20y +150/10 x=−2y+15 solve y 10x+20y+−10x=150+−10x 20y=−10x+150 20y/20=−10x+150/20 y=−1/2 x+15/2 y= 0.5x + 7.5 Does this help?